OpenStudy (anonymous):

a student wrote that -5^0=1 what was his error?

zepdrix (zepdrix):

\[-5^0=-(5)^0\neq (-5)^0\] With the way it was written, the negative was not included in the exponential! :O There should be a - in the answer. Does that make sense? c:

OpenStudy (anonymous):

ya thanks

OpenStudy (anonymous):

order 'o operations \[5^0=1\] \[-5^0=-1\]